OpenStudy (anonymous):

Using the Distance formula, how would you solve from these two points? What do you do with the radicals? (√2,-7√3), (4√2, 8√3)

OpenStudy (anonymous):

just plug them in and solve

OpenStudy (anonymous):

i get √707 but that's not right

OpenStudy (anonymous):

change in x √2 - 4√2 = -4√2 change in y -7√3 - 8√3 = -15√3 then square them change in x: 32 change in y: 675 add them together you get 707

OpenStudy (anonymous):

try \[3\sqrt{77}\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

from \[\sqrt{2}\] to \[4\sqrt{2}\] is \[3\sqrt{2}\] and from \[-7\sqrt{3}\] to \[8\sqrt{3}\] is \[15\sqrt{3}\]

OpenStudy (anonymous):

square. add them together, and then take the square root

OpenStudy (anonymous):

oh i subtracted wrong

OpenStudy (anonymous):

got it. good luck

OpenStudy (anonymous):

thanks

OpenStudy (mertsj):